# supSupplementary References

### v1

@@ Appendix D Task 3 Synthetic Option Generation Procedure and Prompt
We first generate three detailed synthetic options that are specifically crafted to imitate the support judgments from other papers concerning the same type of bias, ensuring that they are tailored to be relevant to the specific paper in question while maintaining the foundational reasoning. See Figure 10 for the prompt template. Following this, we condense these detailed options into shorter versions that preserve the original meaning, using prompt 11. To prevent heuristic algorithms from solving the task easily, we randomly select either the long or short version of each synthetic option to include in the multiple choice questions.
@@
For selecting the three options derived from other papers’ support judgments within the same bias category, we use prompt 12.
@@
Finally, once the six incorrect options are constructed, we conduct a manual review with the help of prompt 13 on all data points. This review ensures that the options are incorrect and not false negatives.
@@
Appendix E Experiment Details
Due to budget limit, we only evaluate the LLMs on a subset of our test set. We provide this subset in our GitHub.
E.1 Prompt Optimization
All evaluation prompts are optimized using disjoint development set for each task. Each prompt includes specialized instructions designed to elicit chain-of-thought reasoning, thereby enhancing the models’ reasoning abilities. These instructions are repeated twice: once at the beginning and once at the end of the prompt, ensuring that the models retain the instructions even after processing the entire paper. Empirical evidence shows that few-shot in-context learning does not improve model performance. Consequently, evaluation prompts do not incorporate in-context learning. It is important to note that prompt optimization is not tailored to any specific large language model (LLM); instead, the aggregated performance across all models is used to guide the optimization of evaluation prompts.
E.2 Prompts for Evaluation
E.2.1 Prompt to Evaluate Study Inclusion/Exclusion (Task 1)
See Figure 14 for prompt to evaluate Study Inclusion/Exclusion (Task 1).
Figure 14: Prompt for Study Inclusion/Exclusion (Task1) evaluation
E.2.2 Prompt to Evaluate Bias Retrieval (Task 2)
See Figure 15 for prompt to evaluate Bias Retrieval (Task 2). We use an additional multi-turn prompt 16 when the model outputs more than required.
Figure 15: Prompt for Bias Retrieval (Task 2) evaluation
Figure 16: Prompt for Bias Retrieval (Task 2) evaluation when the model output more than required
E.2.3 Prompt to Evaluate Support Judgment Selection (Task 3)
See Figure 17 for prompt to evaluate Support Judgment Selection (Task 3).
Figure 17: Prompt for Support Judgment Selection (Task 3) evaluation
E.2.4 Prompt to Evaluate Risk Level Determination (Task 4)
See Figure 18 for prompt to evaluate Risk Level Determination (Task 4).
Figure 18: Prompt for Risk Level Determination (Task 4) evaluation
E.3 Results with Standard Error
We provide bootstrapped standard errors for the experimental results 11,12,13,14.
Table 11: Results for Task 1: Study Inclusion/Exclusion Standard Error Included
F1 Accuracy Weighted Accuracy
Claude3-Opus 75.5 $\pm$ 2.6 77.1 $\pm$ 2.5 74.9 $\pm$ 2.5
Gemini-1.5-Pro 72.2 $\pm$ 2.6 72.7 $\pm$ 2.6 73.0 $\pm$ 2.7
GPT-4o 75.5 $\pm$ 2.6 75.6 $\pm$ 2.5 77.6 $\pm$ 2.3
Table 12: Results for Task 2: Sentence Retrieval Standard Error Included
@@ E.3 Results with Standard Error
Model Full Selection Attrition Performance Detection Reporting Analysis Special Other
n = 337 n = 117 n = 57 n= 50 n = 50 n = 35 n = 13 n = 7 n = 23
Bias Recall@Optimal
Claude3-Opus 39.0 $\pm$ 2.4 55.4 $\pm$ 4.0 22.0 $\pm$ 4.7 57.3 $\pm$ 6.39 46.4 $\pm$ 6.5 11.0 $\pm$ 4.9 47.9 $\pm$ 12.3 13.9 $\pm$ 8.3 4.4 $\pm$ 2.9
Gemini-1.5-Pro 41.1 $\pm$ 2.5 57.2 $\pm$ 4.0 28.7 $\pm$ 5.2 49.1 $\pm$ 6.4 48.8 $\pm$ 6.3 23.3 $\pm$ 6.4 54.2 $\pm$ 11.4 13.9 $\pm$ 8.3 3.3 $\pm$ 2.3
GPT-4o 44.8 $\pm$ 2.5 60.3 $\pm$ 4.0 28.6 $\pm$ 5.0 64.6 $\pm$ 5.5 48.7 $\pm$ 6.4 32.9 $\pm$ 7.3 28.7 $\pm$ 9.7 28.0 $\pm$ 14.1 2.3 $\pm$ 2.2
Bias Recall@3
Claude3-Opus 52.3 $\pm$ 2.5 67.9 $\pm$ 3.8 40.2 $\pm$ 5.8 63.1 $\pm$ 5.9 55.9 $\pm$ 6.6 25.5 $\pm$ 7.1 55.9 $\pm$ 12.3 64.6 $\pm$ 16.9 19.7 $\pm$ 7.8
Gemini-1.5-Pro 56.1 $\pm$ 2.4 69.7 $\pm$ 3.8 44.1 $\pm$ 5.7 73.7 $\pm$ 5.3 65.3 $\pm$ 6.1 35.6 $\pm$ 7.6 56.3 $\pm$ 12.8 35.5 $\pm$ 16.1 6.7 $\pm$ 4.8
GPT-4o 61.8 $\pm$ 2.4 72.2 $\pm$ 3.8 47.9 $\pm$ 5.8 82.0 $\pm$ 4.5 78.5 $\pm$ 5.4 43.2 $\pm$ 8.2 67.9 $\pm$ 10.3 50.4 $\pm$ 17.1 10.3 $\pm$ 5.8
Table 13: Results for Task 3: Support Judgment Selection Standard Error Included
@@
Model Full Selection Attrition Performance Detection Reporting Analysis Special
n = 310 n = 86 n = 60 n = 52 n = 56 n = 50 n = 8 n = 6
Accuracy
Claude3-Opus 51.2 $\pm$ 2.8 62.6 $\pm$ 5.2 56.7 $\pm$ 6.4 48.0 $\pm$ 6.9 46.5 $\pm$ 6.9 44.2 $\pm$ 7.0 50.0 $\pm$ 17.9 0.0 $\pm$ 0.0
Gemini-1.5-Pro 49.3 $\pm$ 3.0 57.1 $\pm$ 5.3 59.9 $\pm$ 6.5 46.4 $\pm$ 7.0 42.8 $\pm$ 6.8 40.2 $\pm$ 7.0 49.3 $\pm$ 17.7 16.6 $\pm$ 15.5
GPT-4o 45.8 $\pm$ 2.9 52.3 $\pm$ 5.4 63.1 $\pm$ 6.2 42.5 $\pm$ 6.7 37.7 $\pm$ 6.5 28.0 $\pm$ 6.2 74.5 $\pm$ 15.7 0.0 $\pm$ 0.0
Table 14: Results for Task 4: Risk Level Determination Standard Error Included
@@
Model Full Selection Attrition Performance Detection Reporting Analysis Special Other
n = 337 n = 93 n = 50 n = 50 n = 50 n = 50 n = 18 n = 26 n = 50
F1
Claude3-Opus 47.7 $\pm$ 2.9 54.3 $\pm$ 6.6 34.1 $\pm$ 7.0 46.7 $\pm$ 6.8 37.4 $\pm$ 7.6 23.0 $\pm$ 4.7 69.9 $\pm$ 18.8 27.9 $\pm$ 6.7 27.9 $\pm$ 1.8
Gemini-1.5-Pro 43.1 $\pm$ 2.6 53.9 $\pm$ 5.8 26.0 $\pm$ 3.7 43.7 $\pm$ 7.2 46.8 $\pm$ 7.4 30.4 $\pm$ 5.2 44.6 $\pm$ 7.7 20.1 $\pm$ 3.4 22.8 $\pm$ 6.1
GPT-4o 47.4 $\pm$ 2.9 54.3 $\pm$ 6.6 30.4 $\pm$ 5.8 46.7 $\pm$ 6.8 37.4 $\pm$ 7.6 22.6 $\pm$ 3.9 69.9 $\pm$ 18.8 32.2 $\pm$ 8.5 27.0 $\pm$ 1.6
Accuracy
Claude3-Opus 56.8 $\pm$ 2.6 64.6 $\pm$ 5.0 58.1 $\pm$ 7.1 48.2 $\pm$ 7.1 41.7 $\pm$ 7.1 40.0 $\pm$ 7.3 88.9 $\pm$ 7.3 53.8 $\pm$ 9.7 72.2 $\pm$ 6.4
Gemini-1.5-Pro 49.3 $\pm$ 2.7 62.5 $\pm$ 5.1 42.4 $\pm$ 7.2 46.5 $\pm$ 6.8 48.1 $\pm$ 7.1 56.0 $\pm$ 7.2 77.5 $\pm$ 9.5 42.4 $\pm$ 9.9 23.7 $\pm$ 6.2
GPT-4o 57.0 $\pm$ 2.6 64.6 $\pm$ 5.0 58.1 $\pm$ 7.2 48.2 $\pm$ 7.1 41.7 $\pm$ 7.2 43.6 $\pm$ 7.2 88.9 $\pm$ 7.3 57.7 $\pm$ 9.7 68.3 $\pm$ 6.5
Weighted Accuracy
Claude3-Opus 49.5 $\pm$ 3.0 55.3 $\pm$ 5.1 36.2 $\pm$ 9.0 48.0 $\pm$ 7.0 46.5 $\pm$ 5.9 30.4 $\pm$ 10.1 69.3 $\pm$ 19.6 28.2 $\pm$ 4.9 32.4 $\pm$ 0.9
Gemini-1.5-Pro 47.9 $\pm$ 3.2 56.3 $\pm$ 5.6 34.0 $\pm$ 10.1 44.4 $\pm$ 7.0 54.2 $\pm$ 8.2 36.4 $\pm$ 10.1 43.6 $\pm$ 4.2 22.2 $\pm$ 4.9 32.5 $\pm$ 10.1
GPT-4o 49.3 $\pm$ 3.1 55.3 $\pm$ 5.1 29.3 $\pm$ 3.9 48.0 $\pm$ 7.0 46.5 $\pm$ 5.9 24.5 $\pm$ 8.4 69.3 $\pm$ 19.6 60.4 $\pm$ 24.4 30.6 $\pm$ 1.5
Appendix F Full Example
We provide the full version of the example presented in the main paper. The example is shown in its entirety here, whereas in the main paper it was shortened to save space.
F.1 Search Protocol Gonzalez et al. [2015]
Types of Studies Randomized controlled trials (RCTs)
Types of Participants People living in leishmaniasis endemic regions.
Types of Interventions Any intervention that aims to reduce leishmaniasis incidence through vector or reservoir control.
Types of Outcome Measures People developing CL or VL infections.
Estimates of the vector density measured by an appropriate technique (adult sandfly density estimated by counts of vectors either landing on exposed body parts of humans acting as baits or collected resting inside buildings, for example, on walls).
Number of participants with positive immunological or biochemical tests that detect contact with the parasite (for example, leishmanin skin test conversion rates or lymphocyte proliferation rates, or both).
Adverse effects on people.
Adherence to control measures; for example, the extent to which specified intervention components were delivered as prescribed.
Measures of environmental impact (assessment of the possible impact - positive or negative - that the interventions may have on the natural environment) or sustainability (assessment of the ability to change biological and human processes, functions, biodiversity and productivity), or both.
F.2 Objective of the Meta-Analysis Gonzalez et al. [2015]
To assess the effects of vector and reservoir control interventions for cutaneous and for visceral leishmaniasis.
F.3 Full Paper
See Werneck et al. for full paper content.
F.4 Bias Name and Definition Higgins and Green [2011]
Bias Name
selective reporting (reporting bias)
Bias Definition
Reporting bias due to selective outcome reporting.
F.5 PICO
Methods Trial design: Cluster-RCT. Unit of randomization: Geographic area. Number of clusters: 40 geographic areas. Entomological data collection: Not done. Clinical data collection: Conversion of the Montenegro skin test (MST) at 18 months of follow-up. Length of follow-up: 18 months. Analysis: Analysed at cluster level.
Participants Ten localities in 7 neighbourhoods of the city of Teresina (Brazil) were divided into blocks, each containing an average of 60 residences. For each locality, 4 blocks were selected to minimize the risk of cross-contamination of interventions. Eligible participants were residents of selected blocks aged 1 year or above with no history of VL. The 40 geographic areas (blocks) randomly allocated to the 4 types of interventions (697 subjects MST-). Endemic disease: VL caused by L. chagasi (L. infantum).
Interventions 1. Spraying households and residential annexes with insecticide. 2. Elimination of infected dogs. 3. Combination of spraying and eliminating infected dogs. 4. No intervention. Description of spraying: performed according to the routine of the VL Control Program of the Zoonosis Control Center of the Teresina City Health Department. Interventions were delivered in the selected blocks every 6 months, for three times, beginning just after each household visit. The elimination of infected dogs was decided if indirect immunofluorescence test was more or equalled 1:40.
Outcomes Cases of infection by L. infantum at 18 months determined by conversion of the MST (MST- at the beginning) or diagnosis of active VL.
Notes Country: Brazil (Teresina, Itararé quarter). Trial dates: January 2004 to December 2006. Trial sponsor: Funded by Health Surveillance Unit from the Brazilian Ministry of Health. One author was partially funded by the Brazilian Research Council (CNPq 306267/2010-1 and 202088/2012-0). The founders had no role in trial design, data collection and analysis, decision to publish, or preparation of the manuscript. The authors have declared that no competing interests exist. Sample size: Calculated. Compliance assessment: Not reported.
[…]

### v2

@@ Appendix H SJS Synthetic Option Generation Procedure and Prompt
We first generate three detailed synthetic options that are specifically crafted to imitate the support judgments from other papers concerning the same type of bias, ensuring that they are tailored to be relevant to the specific paper in question while maintaining the foundational reasoning. See Figure 11 for the prompt template. Following this, we condense these detailed options into shorter versions that preserve the original meaning, using prompt 12. To prevent heuristic algorithms from solving the task easily, we randomly select either the long or short version of each synthetic option to include in the multiple choice questions.
@@
For selecting the three options derived from other papers’ support judgments within the same bias category, we use prompt 13.
@@
Finally, once the six incorrect options are constructed, we conduct a manual review with the help of prompt 14 on all data points. This review ensures that the options are incorrect and not false negatives.
@@
Appendix I Experiment Details
I.1 Prompt Optimization
All evaluation prompts are optimized using disjoint development set for each task. Each prompt includes specialized instructions designed to elicit chain-of-thought reasoning, thereby enhancing the models’ reasoning abilities. These instructions are repeated twice: once at the beginning and once at the end of the prompt, ensuring that the models retain the instructions even after processing the entire paper. Empirical evidence shows that few-shot in-context learning does not improve model performance. Consequently, evaluation prompts do not incorporate in-context learning. It is important to note that prompt optimization is not tailored to any specific large language model; instead, the aggregated performance across all models is used to guide the optimization of evaluation prompts.
I.2 Prompts for Evaluation
I.2.1 Prompt to Evaluate Support Sentence Retrieval (SSR)
See Figure 15 for prompt to evaluate Support Sentence Retrieval (SSR). We use an additional multi-turn prompt 16 when the model outputs more than required.
Figure 15: Prompt for Support Sentence Retrieval (SSR) evaluation.
Figure 16: Prompt for Support Sentence Retrieval (SSR) evaluation when the model output more than required.
I.2.2 Prompt to Evaluate Support Judgment Selection (SJS)
See Figure 17 for prompt to evaluate Support Judgment Selection (SJS).
Figure 17: Prompt for Support Judgment Selection (SJS) evaluation.
I.2.3 Prompt to Evaluate Risk-of-Bias Determination (Main Task)
See Figure 18 for prompt to evaluate Risk-of-Bias Determination (Main Task).
Figure 18: Prompt for Risk-of-Bias Determination (Main Task) evaluation.
I.3 Prompt for Risk-of-Bias Determination (Main Task) Given Retrieval Results
See Figure 19 for prompt for Risk-of-Bias Determination (Main Task) given retrieval results.
Figure 19: Prompt for Risk-of-Bias Determination (Main Task) Given Retrieval Results.
I.4 Prompt for OpenAI o4-mini to identify support sentences
See Figure 20 for prompt for OpenAI o4-mini to identify support sentences
Figure 20: Prompt for OpenAI o4-mini to identify support sentences.
I.5 Results with Standard Error
We provide bootstrapped standard errors for the experimental results in Tables 13,14,15.
@@ I.5 Results with Standard Error
Model Avg Selection Attrition Performance Detection Reporting
n = 157 n = 46 n= 54 n = 61 n = 14
OpenAI-v3 22.72 40.13 $\pm$ 3.56 6.3 $\pm$ 2.72 26.67 $\pm$ 5.29 27.38 $\pm$ 4.81 13.1 $\pm$ 7.56
GritLM-7B 18.87 35.84 $\pm$ 3.59 6.99 $\pm$ 2.67 15.19 $\pm$ 3.76 26.83 $\pm$ 5.08 9.52 $\pm$ 7.09
GPT-4o 47.48 60.5 $\pm$ 3.54 42.36 $\pm$ 6.47 41.48 $\pm$ 6.06 50.05 $\pm$ 5.7 43.03 $\pm$ 11.55
Sonnet-3.5 39.17 55.42 $\pm$ 3.62 30.65 $\pm$ 5.71 37.01 $\pm$ 5.69 37.21 $\pm$ 5.5 35.54 $\pm$ 11.55
Llama-3.1-70B 45.62 61.17 $\pm$ 3.48 34.21 $\pm$ 5.95 45.19 $\pm$ 5.96 50.11 $\pm$ 5.68 37.41 $\pm$ 11.49
Llama-3-8B 22.66 49.44 $\pm$ 3.68 14.49 $\pm$ 4.86 22.13 $\pm$ 4.95 20.77 $\pm$ 4.7 6.46 $\pm$ 4.4
Llama-3-8B Fine-Tuned 40.8 69.01 $\pm$ 3.41 24.78 $\pm$ 5.85 47.28 $\pm$ 6.26 48.63 $\pm$ 5.92 14.29 $\pm$ 9.35
Table 13: Results for Sentence Retrieval Standard Error Included.
@@
Model Full Selection Attrition Performance Detection Reporting
n = 310 n = 86 n = 60 n = 52 n = 56 n = 50
GPT-4o 47.17 58.46 $\pm$ 4.37 60.2 $\pm$ 5.04 48.28 $\pm$ 5.24 42.68 $\pm$ 5.29 26.25 $\pm$ 4.7
Sonnet-3.5 59.92 73.08 $\pm$ 3.89 73.47 $\pm$ 4.28 50.57 $\pm$ 5.42 51.22 $\pm$ 5.44 51.25 $\pm$ 5.26
Llama-3.1-70B 53.16 66.15 $\pm$ 3.89 62.24 $\pm$ 4.98 44.83 $\pm$ 5.19 46.34 $\pm$ 5.64 46.25 $\pm$ 5.7
Llama-3-8B 26.54 26.92 $\pm$ 3.87 34.69 $\pm$ 4.77 24.14 $\pm$ 4.64 21.95 $\pm$ 4.69 25.0 $\pm$ 4.91
Llama-3-8B Fine-Tuned 29.6 40.77 $\pm$ 4.36 28.57 $\pm$ 4.37 24.14 $\pm$ 4.42 19.51 $\pm$ 4.52 35.0 $\pm$ 5.38
Table 14: Results for Support Judgment Selection Standard Error Included.
@@
Model Avg Selection Attrition Performance Detection Reporting Deviation
n = 933 n = 629 n = 309 n = 645 n = 594 n= 331
GPT-4o 42.07 50.52 $\pm$ 1.87 36.48 $\pm$ 2.14 54.67 $\pm$ 2.8 43.66 $\pm$ 2.02 34.43 $\pm$ 1.87 32.65 $\pm$ 2.77
Sonnet-3.5 41.93 52.83 $\pm$ 1.86 37.07 $\pm$ 2.04 50.56 $\pm$ 2.75 43.29 $\pm$ 2.09 33.58 $\pm$ 1.41 34.23 $\pm$ 1.96
Llama-3.1-70B 38.81 48.1 $\pm$ 1.85 31.7 $\pm$ 1.94 48.04 $\pm$ 2.67 44.36 $\pm$ 2.0 31.58 $\pm$ 1.33 29.04 $\pm$ 2.37
Llama-3-8B 30.05 36.37 $\pm$ 1.28 32.44 $\pm$ 1.62 39.13 $\pm$ 2.59 37.21 $\pm$ 1.92 19.76 $\pm$ 1.96 15.38 $\pm$ 1.64
Llama-3-8B Fine-Tuned 36.33 49.51 $\pm$ 1.91 33.34 $\pm$ 1.89 42.44 $\pm$ 2.69 40.1 $\pm$ 2.01 26.95 $\pm$ 0.44 25.67 $\pm$ 1.06
Table 15: Results for Risk Level Determination Standard Error Included.